OpenStudy (anonymous):

Fill in the blank Cos _____ equals √3/2 A. 75 B. 30 C. 45 D. 60

OpenStudy (anonymous):

find inverse cos (sqrt(3)/2)

OpenStudy (anonymous):

How do I do that?

OpenStudy (anonymous):

ok, one moment. I'm close.

OpenStudy (anonymous):

ok, imagine you have a 30 60 90 triangle  cos(30) = sqrt(3)/2 inverse cos (30) = 1/cos(30) 1/cos(30) = 1/(sqrt(3)/2)) 1/cos(30) = 2/(sqrt(3))